# From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks

### v1

@@
Abstract
Compositionality has long been considered a key explanatory property underlying human intelligence: arbitrary concepts can be composed into novel complex combinations, permitting the acquisition of an open ended, potentially infinite expressive capacity from finite learning experiences. Influential arguments have held that neural networks fail to explain this aspect of behavior, leading many to dismiss them as viable models of human cognition. Over the last decade, however, modern deep neural networks (DNNs)—which share the same fundamental design principles as their predecessors—have come to dominate artificial intelligence, exhibiting the most advanced cognitive behaviors ever demonstrated in machines. In particular, large language models (LLMs), DNNs trained to predict the next word on a large corpus of text, have proven capable of sophisticated behaviors such as writing syntactically complex sentences without grammatical errors, producing cogent chains of reasoning, and even writing original computer programs—all behaviors thought to require compositional processing. In this chapter, we survey recent empirical work from machine learning for a broad audience in philosophy, cognitive science, and neuroscience, situating recent breakthroughs in deep learning and LLMs within the broader context of philosophical arguments about compositionality. In particular, our review emphasizes two approaches to endowing neural networks with compositional generalization capabilities: (1) architectural inductive biases, and (2) metalearning, or learning to learn. We also present findings suggesting that LLM pretraining can be understood as a kind of metalearning, and can thereby equip DNNs with compositional generalization abilities in a similar way. We conclude by discussing the implications that these findings may have for the study of compositionality in human cognition and by suggesting avenues for future research.
@@ 1 Introduction
Deep neural networks (DNNs) have made breakthrough after breakthrough in artificial intelligence (AI) over the last decade, reproducing sophisticated cognitive behaviors from advanced gameplay in board games like Go (Silver et al.,, 2016), Chess (Silver et al.,, 2018), and Diplomacy (FAIR et al.,, 2022), to generating stunning images from arbitrary natural language instructions (Betker et al.,, 2023), to catalyzing achievements in mathematics (Romera-Paredes et al.,, 2024), science (Jumper et al.,, 2021), engineering (Merchant et al.,, 2023), and medicine (Omiye et al.,, 2023; Singhal et al.,, 2023). The large language models (LLMs) powering AI products like chatGPT (Brown et al.,, 2020; Bubeck et al.,, 2023; OpenAI et al.,, 2024) have shown remarkable proficiency in cognitive domains previously regarded as uniquely human, such as natural language syntax (Linzen and Baroni,, 2021), critical reasoning and argumentation (Herbold et al.,, 2023), and computer programming (Chen et al.,, 2021; Bubeck et al.,, 2023).
@@
In humans, these creative cognitive behaviors have traditionally been explained in part by postulating a special cognitive property: compositionality. In particular, the open-ended systematic and productive generalization behaviors exhibited by humans are often modeled by the compositionality of classical cognitive architectures, which intrinsically guarantees that they can redeploy familiar constituents in novel constructions (Fodor and Pylyshyn,, 1988). Philosophers and cognitive scientists have argued that compositionality should be considered non-negotiable in this respect, the “only game in town” for explaining the power and flexibility of human intelligence across a number of domains, particularly natural language syntax and semantics (Fodor,, 1975, 2000).
@@
Unlike classical symbolic AI systems, however, the neural networks powering these recent advances are not designed to embody a strong assumption of compositionality, relying instead on acquiring capabilities by learning statistical regularities from data (Fodor and Pylyshyn,, 1988; Smolensky et al., 2022a, ; McClelland et al., 1986b, ; Rumelhart et al., 1986b, ). Whereas compositionality is integral to certain fundamental design features of classical computational architectures (e.g., predicate-argument structure, variable binding, constituent structure), these features are not native to standard DNNs. Traditionally, this has been lodged as an objection to the viability of neural networks as cognitive models (Fodor and Pylyshyn,, 1988). However, the advances in new DNNs’ capabilities in domains traditionally thought to require compositionality (e.g., language) calls this into question. Have today’s models overcome previously hypothesized limitations, or are they merely imitating compositional behaviors through useful but imperfect heuristics? Does the fact that DNNs reproduce sophisticated cognitive behaviors demonstrate that there is a new game in town and that compositionality is negotiable after all? Or have modern DNNs achieved genuine compositionality by non-classical means?
@@
In this chapter, we’ll review and contextualize recent studies of compositionality in DNNs, with a particular focus on language. We begin in section 2.1 with a historical overview of the explanatory role of compositionality in philosophy, linguistics, and psychology, outlining its evolution from a property of interest primarily to logicians and philosophers to an empirically motivated psychological construct central to contemporary cognitive science. We then briefly describe the rise of connectionism in the 1980s and 90s (section 2.2) and the challenge that compositionality poses for neural networks (section 2.3). With this crucial context in place, we move to today’s deep neural networks, providing a short overview of recent advances, with an emphasis on the transformer architecture (section 3). We review key recent breakthroughs in replicating compositional behaviors in DNNs (section 4), and argue that architectural inductive biases (section 4.2) and metalearning (sections 4.3 and 4.4) offer novel approaches to reproducing the behavioral signatures of compositionality in neural networks. To close, we discuss potential implications that this work may have for our understanding of human compositionality, and suggest avenues for future empirical and philosophical work (section 5).
@@ 2.1 From Frege to Fodor
Though the compositionality of natural language and thought is a central commitment of contemporary cognitive science, its current explanatory role was not always evident. Appreciation of the significance of compositionality first emerged at a time when philosophy, psychology, logic, and linguistics were still intertwined in an undifferentiated study of the mind (Pagin and Westerståhl,, 2010), and it took nearly a century of philosophical investigation to cleanly excavate the question(s) that the compositionality principle purports to answer. [^1] Compositionality can be traced even further back in the history of philosophy, to Aristotle and philosophers in the Indian tradition (Pagin and Westerståhl,, 2010), but our focus here will be on the analytic tradition, within which this notion achieved its current position of prominence and its very name (Putnam,, 1995; Katz and Fodor,, 1963).
@@
Gottlob Frege made indispensable contributions to the disentangling of these domains (Dummett,, 1993) and was arguably the first philosopher to appreciate the explanatory potential of the compositionality principle. In contemporary terms, the compositionality principle states that the meaning of a complex expression is determined by the meaning of its parts and the way they are combined. There are a number of different ways of precisifying this principle, [^2] For discussion, see Pagin and Westerståhl, (2010) and Partee, (2004). and two importantly different ways of interpreting its significance—either as a claim about meaning itself or as a claim about the psychological processes involved in generating and grasping meanings. On the first interpretation, the compositionality principle asserts that the meaning of a complex expression is exhaustively determined by (or supervenes on) the meanings of its parts and their mode of combination. [^3] On this reading, the compositionality principle is a claim about the metaphysics of meaning, akin to the mind-brain supervenience thesis, which asserts that one’s psychological states are determined by their physical or neural states (Kim,, 1981). This first sense may be less familiar to a cognitive science and machine learning audience, but it has played a key role in long-standing debates about compositionality within philosophy. On the second, it expresses a psychological hypothesis about how language users themselves determine this meaning—namely, by assembling the meanings of the parts in accordance with tacitly known rules (Dummett,, 1996).
@@
Though Frege clearly appreciated the significance of compositionality and anticipated key contemporary arguments in its favor (Frege,, 1956, 1980), it is not clear which of these two construals of the compositionality principle, if any, he intended. In fact, though it been dubbed “Frege’s Principle” (Cresswell,, 1973), Frege never explicitly formulated the compositionality principle himself. The first philosopher to do so was Frege’s one-time student, Rudolf Carnap, who named it after his former teacher (Carnap,, 1988). [^4] There is significant scholarly dispute over Frege’s views on the compositionality principle (Pelletier,, 2001; Janssen,, 2012). Janssen suggests that it would better be called “Carnap’s Principle.”
@@
Carnap was the leading philosopher of the logical empiricism movement and, among other contributions, paved the way for what was arguably the central research program in 20th century formal semantics (Carnap,, 1988; Montague,, 1974). Carnap employed "Frege’s Principle" as a crucial theoretical assumption. However, like Frege before him, Carnap was deeply skeptical that these formal languages could be treated as models of natural (“vernacular”) languages (like German, English, or Mandarin) (Carnap,, 1937). His theoretical ambition was not to explain everyday language use or uncover underlying psychological mechanisms, as a contemporary theorist might expect, but to develop artificial languages ideally suited to mathematical and scientific purposes (Harris,, 2017). Compositionality was an indispensable feature of these artificial languages, but Carnap was well aware of the idealizations that they involved—abstracting away from central features of natural language, like context-sensitivity and ambiguity—and self-consciously avoided putting them forward as models of natural language. [^5] This is why Harris, (2017) frames Carnap and other ‘ideal language philosophers’ as part of the “pre-history” of natural language semantics, rather than its history proper. This held for philosophers working in the ‘ideal language tradition’ more broadly. They appreciated the power of mathematical logic and the importance of the compositionality principle, but did not advance their formal theories as psychological models or treat compositionality as an empirical hypothesis about human linguistic processing.
@@
Ordinary Language Philosophy, the other major branch of 20th century analytic philosophy’s ‘linguistic turn’ (Rorty,, 1992), criticized the ideal language philosophers for this abstraction and focused their analyses on the messy realities of natural language use (Ryle,, 1949). Leaving aside the movement’s grander ambitions (Hacker,, 1975; Horwich,, 2013), philosophers in the ordinary language tradition made clear-cut contributions to linguistic theorizing, from J.L. Austin’s work on speech acts (Austin,, 1975) to H.P. Grice’s pioneering contributions to the field of pragmatics (Grice,, 1991). However, these philosophers were either skeptical or silent about the compositionality of natural language. To take an illustrative example, Gilbert Ryle, a leading ordinary language philosopher and philosophical behaviorist, rejected the compositionality principle and the ‘building block’ theory of meaning in which it was at home (Ryle,, 1957). Following Wittgenstein, (1973), Ryle argued that linguistic meaning is determined by use (Ryle,, 1957, 1964), [^6] This Wittgensteinian perspective on linguistic meaning inspired J.R. Firth’s distributional semantics (Firth,, 1957), a key historical antecedent of the statistical approaches powering today’s large language models (Brunila and LaViolette,, 2022). and that use is highly contextual—trying to extract a single, unified contribution that a given word makes to each of the sentences (or speech acts) in which it may occur would be a methodological dead-end. Put simply, in Ryle’s view, the use theory of meaning runs counter to compositionality, so the compositionality principle should be rejected.
@@
The tension between these approaches to the study of language persisted into the early 1960s, with ideal language philosophers developing compositional models, but abstracting away from natural language use, and ordinary language philosophers advancing analyses of natural language, but rejecting or simply ignoring the principle of compositionality. This contrast formed the philosophical backdrop for Noam Chomsky’s revolution in the field of syntax (Chomsky,, 1957, 1965), which would fundamentally alter the role of compositionality in linguistic theorizing.
@@
Chomsky rejected the key assumption that both groups of philosophers shared—that formal mathematical theories could not be treated as models of ordinary language. This allowed him to synthesize their approaches, harnessing the “methodological toolkit” developed by philosophers interested in formal logic, from Frege on (Harris,, 2017), but taking them seriously as models of the syntax of natural languages (Chomsky,, 1957; Katz,, 1971). Though Chomsky did not focus on questions of meaning, and would eventually express some skepticism regarding the compositionality of natural languages (Chomsky,, 1975), the synthesis that his work enacted was a crucial step in the emergence of the compositionality principle as an empirical hypothesis about human linguistic processing. Indeed, it was promptly extended to formal semantics by Richard Montague and other key contributors (Katz and Fodor,, 1963; Davidson,, 1965). Montague explicitly argued that there is “no important theoretical difference between natural languages and the artificial languages of logicians,” allowing him to treat rule-based semantic systems of the kind pioneered by Carnap as models of human linguistic competence (Montague,, 1974, p. 15). [^7] Montague signals his alignment with Chomsky on this key point: “I consider it possible to comprehend the syntax and semantics of both kinds of languages within a single natural and mathematically precise theory. On this point I differ from a number of philosophers, but agree, I believe, with Chomsky and his associates” (Montague,, 1974, p.15).
@@
It would be hard to overstate the impact that this shift had on linguistic theorizing. In effect, it transformed formal syntactic and semantic theories from abstract objects of study for logicians, mathematicians, and philosophers into psychological constructs that underlie and explain the linguistic capacities of even the youngest children capable of speech.
@@
With this shift, the compositionality principle took on a newfound significance and key role in explaining human linguistic behavior. Chomsky drew attention to the creativity of human language use and its open-ended, potentially infinite expressive capacity, which must be achieved through finite means (Chomsky,, 1966, 2020; Frege,, 1980). To use a classic example discussed by Katz and Fodor, (1963), even in the most routine linguistic interactions we regularly encounter novel sentences, yet fluidly and effortlessly comprehend their meaning and generate appropriate responses. This open-ended creative capacity seems to cry out for explanation. [^8] Chomsky suggests that some 17th century philosophers, like Descartes and Wilhelm von Humboldt, recognized this explanatory challenge and anticipated some of his own ideas (Chomsky,, 1966). It is also prefigured by Frege, who wrote in 1914 that, “the possibility of our understanding sentences which we have never heard before rests evidently on this, that we can construct the sense of a sentence out of parts that correspond to words… Without this, language in the proper sense would be impossible… we would always be restricted to a very narrow area and could not form a completely new proposition” (Frege,, 1980). How are humans capable of understanding sentences that they have never before encountered?
@@
The answer, in one word, is compositionality. If the meaning of an unfamiliar sentence is determined by the meaning of its parts, their mode of combination, and nothing else, then knowing the meanings of component words and the grammar of the language will suffice for grasping this novel meaning. Appealing to tacit knowledge (Fodor, 1968a, ; Evans,, 1985) thus provides an explanation of the human capacity to understand novel sentences—we derive the meaning of novel constructions by recombining known components according to known syntactic rules.
@@
In fact, it can seem that tacit knowledge of a compositional grammar is the only adequate explanation of this capability. Chomsky argued that humans’ creative capacity is essentially unbounded (Chomsky,, 1957). To take a classic example, if we idealize away from constraints of memory and attention (as well as the finitude of a human lifetime!), a compositional subject could go on adding new conjuncts to the end of a well-understood sentence ad infinitum. Provided that they understand the meaning of the new conjunct (and conjunction itself), they should also be able to understand these newly formed sentences.
It is this open-endedness that makes an internalized compositional system seem like “the only game in town.” Alternative explanatory approaches of the time, like behaviorism, struggled to account for this form of unbounded extension. A behaviorist account of such capacities in terms of simpler psychological mechanisms like stimulus-response association (via classical or operant conditioning), would fail to explain these generalization behaviors involving previously unencountered stimuli. For Chomsky, Montague, Fodor, and others, the clear solution was to posit further internal compositional mechanisms that support these behavioral capacities.
@@
This abductive inference to the best explanation forges a close link between creativity, productivity, and compositionality. Particular observations about productive or systematic behaviors in humans motivate the ascription of internal compositional mechanisms (which behaviorists would never have countenanced) to ordinary speakers. Such arguments commit theorists to interpreting the principle of compositionality as a psychological, rather than a metaphysical, thesis. This point deserves emphasis. To serve as a robust explanation of human language comprehension and production, this compositional system must actually get deployed by humans when they generate the behaviors that the compositionality principle purports to explain. [^9] This observation may seem obvious, but it is in fact easy to overlook. Even Davidson, one of the earliest and most influential proponents of compositional natural language semantics, fails to fully appreciate the point. For a detailed discussion of this tension in Davidson’s work, see Heck, (2013). For the proposed explanation to work, it is not enough that the language itself admits of a compositional analysis—after all, one and the same language may admit both compositional and non-compositional models.
@@
To see the point, consider a simplified ‘toy’ language (Evans,, 1985) with only 10 names (a, b, c…) and 10 predicates (F, G, H…), generating 100 possible sentences (Fa, Fb, Ga…). This language can be provided with a compositional model, along familiar lines. However, every admissible sentence in this language could also be generated by a very different, though ‘extensionally equivalent,’ model that has 100 different axioms specifying the meaning of each. How can we tell which of these models a speaker of this minimal language possesses? Many philosophers, from W. V. O. Quine, (1970) to Crispin Wright, (1981), were skeptical that such questions could be made empirically or behaviorally tractable. Evans, (1985), however, argues that there are clear empirical considerations that justify the ascription of the compositional system rather than the non-compositional one. In principle, we could scrutinize the speaker’s internal states and attempt to detect compositional structure. In practice—and in keeping with the methodological assumptions of researchers in the Chomskyan tradition—we can instead rely on a wider class of behavior—compositional generalization behavior involving the introduction of novel names or predicates (if you introduce a new sentence, Fz, will the speaker be able to understand Gz and Hz?). Evans’ proposal about how to operationalize compositionality is very similar to those used in contemporary machine learning research (see section 4.1), and it demonstrates the central role that linguistic behavior plays in motivating cognitive explanation in terms of the compositionality principle.
@@
These theoretical developments close the gap between early philosophical theorizing about compositionality and the ongoing debates in contemporary cognitive science. Compositionality is understood as an explanatory psychological property, operationalized in terms of the specific kinds of productive and systematic generalization behaviors that humans exhibit (Evans,, 1985). In each case, the characteristic ‘behavioral signatures’ of compositionality are taken as evidence diagnostic of internalized compositional mechanisms, which serve to explain the behaviors in question. As Chomsky often emphasizes, this means that even the most abstract reaches of formal linguistics are properly understood as a branch of cognitive psychology (and ultimately biology, though developing an account of how symbolic systems are implemented in the brain is sometimes treated as a subsidiary enterprise; Chomsky,, 1968).

### v2

@@
From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks
Jacob Russin
Department of Computer Science
Brown University
jake_russin@brown.edu
&Sam Whitman McGrath11footnotemark: 1
Department of Philosophy
Fordham University
smcgrath20@fordham.edu
[…]
Department of Philosophy
Psychological and Brain Sciences
Washington University in St. Louis
Equal primary contribution.
@@ 1 Introduction
Deep neural networks (DNNs) have made breakthrough after breakthrough in artificial intelligence (AI) over the last decade, reproducing sophisticated cognitive behaviors from advanced gameplay in board games like Go (Silver et al.,, 2016), Chess (Silver et al.,, 2018), and Diplomacy (FAIR et al.,, 2022), to catalyzing achievements in mathematics (Romera-Paredes et al.,, 2024), science (Jumper et al.,, 2021), engineering (Merchant et al.,, 2023), and medicine (Omiye et al.,, 2023; Singhal et al.,, 2023). The large language models (LLMs) powering AI products like chatGPT (Brown et al.,, 2020; Bubeck et al.,, 2023; OpenAI et al.,, 2024) have shown remarkable proficiency in cognitive domains previously regarded as uniquely human, such as natural language syntax (Linzen and Baroni,, 2021), critical reasoning and argumentation (Herbold et al.,, 2023), and computer programming (Chen et al.,, 2021; Bubeck et al.,, 2023).
@@
In humans, these creative cognitive behaviors have traditionally been explained by postulating a special cognitive property: compositionality. The precise definition of compositionality varies by discipline, but intuitively it can be understood as the ability to compose familiar constituents (words, concepts, percepts) into novel complex combinations. In particular, cognitive scientists have often modeled human generalization behaviors with classical architectures that intrinsically guarantee that familiar constituents can be redeployed in novel constructions (Fodor and Pylyshyn,, 1988). Theorists have argued that this explanatory framework is ‘non-negotiable,’ that compositionality is the “only game in town” for explaining the power and flexibility of human intelligence across many domains, including natural language syntax and semantics (Fodor,, 1975).
@@
Unlike classical symbolic AI systems, however, the neural networks powering recent advances are not designed to embody a strong assumption of compositionality, relying instead on acquiring capabilities by learning statistical regularities from data (Fodor and Pylyshyn,, 1988; McClelland et al., 1986b, ; Rumelhart et al., 1986b, ; Smolensky et al.,, 2022). Whereas compositionality is integral to certain fundamental design features of classical computational architectures (e.g., predicate-argument structure, variable binding, constituent structure), these features are not native to standard DNNs. Traditionally, this has been lodged as an objection to the viability of neural networks as cognitive models (Fodor and Pylyshyn,, 1988). However, modern DNNs’ impressive performance in domains traditionally thought to require compositionality (e.g., language) calls this into question. Have today’s models overcome previously hypothesized limitations? Or are they merely imitating compositional behaviors through useful but imperfect heuristics? Do DNNs instantiate a novel form of non-classical compositionality? Or does their success in reproducing sophisticated cognitive behaviors show that there is a new game in town and that compositionality is negotiable after all?
@@
In this chapter, we’ll review and contextualize recent studies of compositionality in DNNs, with a particular focus on language. We begin in section 2.1 with a historical overview of the explanatory role of compositionality in philosophy, linguistics, and psychology, outlining its evolution from a property of interest primarily to logicians and philosophers to an empirically motivated psychological construct central to contemporary cognitive science. We then briefly describe the rise of connectionism in the 1980s and 90s (section 2.2) and the challenge that compositionality poses for neural networks (section 2.3). With this crucial context in place, we move to today’s deep neural networks, reviewing key recent breakthroughs in replicating compositional behaviors (section 3). In particular, we argue that metalearning (section 3.2), or learning to learn, offers a novel perspective on how neural networks like LLMs (section 3.3) can come to reproduce the behavioral signatures of compositionality. To close, we discuss potential implications that this work may have for our understanding of human compositionality, and suggest avenues for future empirical and philosophical work (section 4).
@@ 2.1 From Frege to Fodor
Though the compositionality of human language and thought is a central commitment of contemporary cognitive science, its current explanatory role was not always evident. Appreciation of the significance of compositionality first emerged at a time when philosophy, psychology, logic, and linguistics were still intertwined, and it took nearly a century of philosophical investigation to cleanly excavate the question(s) that the compositionality principle purports to answer. [^1] Compositionality can be traced even further back in the history of philosophy, to Aristotle and the Sanskrit grammarians (Pagin and Westerståhl,, 2010), but our focus here will be on the 20th century analytic tradition, in which this phenomenon was first named and analyzed (Putnam,, 1995; Katz and Fodor,, 1963).
@@
Gottlob Frege made indispensable contributions to the disentangling of these domains and was arguably the first philosopher to fully appreciate the explanatory potential of the compositionality principle (Dummett,, 1993). In contemporary terms, the compositionality principle states that the meaning of a complex expression is determined by the meaning of its parts and the way they are combined. There are a number of different ways of precisifying this principle, [^2] For discussion, see Pagin and Westerståhl, (2010) and Partee, (2004). and two importantly different ways of interpreting its significance—as a claim about the metaphysics of meaning or as a claim about the psychological processes with which humans generate and grasp meanings. On the first interpretation, the compositionality principle asserts that the meaning of a complex expression (e.g. a sentence) is determined by or supervenes on the meanings of its parts (e.g. words), which are explanatorily or metaphysically fundamental. On the second, it expresses a psychological hypothesis about how language users themselves determine this meaning—namely, by assembling the meanings of the parts in accordance with tacitly known rules (Dummett,, 1996). [^3] These two readings might be tethered to a distinction between two different senses of “determination.” Consider an analogy. Objectively speaking, the solution to the equation “a x b = ?” is determined by the values of “a” and “b” (as well as the nature of multiplication). Psychologically speaking, however, there are numerous ways that someone might determine this solution—by searching their memory, by guessing, or by actually multiplying the component values. Our question is which of these senses of determination, the objective or the subjective sense (so to speak), is involved in the compositionality principle.
@@
Though Frege clearly appreciated the significance of compositionality and anticipated key contemporary arguments in its favor (Frege,, 1956, 1980), it is not clear which of these two construals of the compositionality principle he intended. In fact, though it has been dubbed “Frege’s Principle” (Cresswell,, 1973), Frege never explicitly formulated the compositionality principle himself. The first philosopher to do so was Frege’s one-time student, Rudolf Carnap, who named it after his former teacher (Carnap,, 1988).
@@
Carnap was the leading proponent of logical empiricism and, among other contributions, paved the way for what was arguably the central research program in 20th century formal semantics (Carnap,, 1988; Montague,, 1974). Carnap employed “Frege’s Principle” as a crucial theoretical assumption. However, like Frege before him, Carnap was deeply skeptical that compositional formal languages could be treated as models of natural languages, like German, English, or Mandarin (Carnap,, 1937). His theoretical ambition was not to explain everyday language use or uncover underlying psychological mechanisms, as a contemporary theorist might expect, but to develop artificial languages ideally suited to mathematical and scientific purposes (Harris,, 2017). Compositionality was an indispensable feature of these artificial languages, but Carnap was well aware of the idealizations that they involved—abstracting away from central features of natural language, like context-sensitivity and ambiguity—and he consciously avoided putting them forward as models of “vernacular” human language. This was true for philosophers in the ‘ideal language tradition’ more broadly. They appreciated the power of mathematical logic and the importance of the compositionality principle, but did not advance their formal theories as psychological models or treat compositionality as an empirical hypothesis about human linguistic processing.
@@
Ordinary Language Philosophy, the other major offshoot of the analytic tradition’s ‘linguistic turn’ (Rorty,, 1992), criticized the ideal language philosophers for this abstraction and focused their analyses on the messy realities of natural language use (Ryle,, 1949). Leaving aside the movement’s grander ambitions (Hacker,, 1975; Horwich,, 2013), philosophers in the ordinary language tradition made clear-cut contributions to linguistic theorizing, from J.L. Austin’s work on speech acts (Austin,, 1975) to H.P. Grice’s pioneering contributions to the field of pragmatics (Grice,, 1991). However, these philosophers were either skeptical or silent about the compositionality of natural language (see, for example, Gilbert Ryle’s attack on the ‘building block’ theory of meaning in Ryle,, 1957).
@@
The tension between these approaches persisted into the early 1960s, with ideal language philosophers developing compositional models, but abstracting away from natural language use, and ordinary language philosophers advancing analyses of natural language, but rejecting or simply ignoring the principle of compositionality. This contrast formed the philosophical backdrop for Noam Chomsky’s revolution in the field of syntax (Chomsky,, 1957, 1965), which would help inaugurate the ‘cognitive turn’ and fundamentally alter the role of compositionality in linguistic theorizing.
@@
Chomsky rejected the key assumption that both groups of philosophers shared—that formal mathematical languages could not be treated as psychological models of ordinary linguistic knowledge. This allowed him to synthesize their approaches, using the “methodological toolkit” developed by philosophers interested in formal logic, from Frege onwards (Harris,, 2017), but taking them seriously as models of the syntax of natural languages (Chomsky,, 1957; Katz,, 1971). Chomsky did not focus on questions of meaning, but the synthesis that his work enacted was a crucial step in the emergence of the compositionality principle as an empirical hypothesis about human linguistic processing. Indeed, Chomsky’s synthesis was promptly extended to formal semantics by Richard Montague and other key contributors (Katz and Fodor,, 1963; Davidson,, 1965). Montague explicitly argued that there is “no important theoretical difference between natural languages and the artificial languages of logicians,” allowing him to treat rule-based semantic systems of the kind pioneered by Carnap as psychological models of human linguistic competence (Montague,, 1974, p. 15). [^4] Montague signals his alignment with Chomsky on this key point: “I consider it possible to comprehend the syntax and semantics of both kinds of languages within a single natural and mathematically precise theory. On this point I differ from a number of philosophers, but agree, I believe, with Chomsky and his associates” (Montague,, 1974, p.15).
@@
It would be hard to overstate the impact that this shift had on linguistic theorizing. In effect, it transformed formal syntactic and semantic theories from abstract objects of study for logicians, mathematicians, and philosophers into psychological constructs that underlie and explain the linguistic capacities of even the youngest children capable of speech. With this shift, the compositionality principle took on a newfound significance and key role in explaining human linguistic behavior.
@@
Chomsky drew attention to the creativity of human language use and its open-ended, potentially infinite expressive capacity, which must be achieved through finite means (Chomsky,, 1966, 2020; Frege,, 1980). To use a classic example discussed by Katz and Fodor, (1963), in even the most routine linguistic interactions we regularly encounter novel sentences, yet fluidly and effortlessly comprehend their meaning and generate appropriate responses. This open-ended creative capacity seems to cry out for explanation. [^5] Chomsky suggests that some 17th century philosophers, like Descartes and Wilhelm von Humboldt, recognized this explanatory challenge and anticipated his own ideas (Chomsky,, 1966). It is also prefigured by Frege, who wrote in 1914 that, “the possibility of our understanding sentences which we have never heard before rests evidently on this, that we can construct the sense of a sentence out of parts that correspond to words… Without this, language in the proper sense would be impossible… we would always be restricted to a very narrow area and could not form a completely new proposition” (Frege,, 1980). How are humans capable of understanding sentences that they have never before encountered?
@@
The answer, in one word, is compositionality. If the meaning of an unfamiliar sentence is determined by the meaning of its parts, their mode of combination, and nothing else, then knowing the meanings of component words and the grammar of the language will suffice for grasping this novel meaning. Appealing to tacit knowledge (Fodor,, 1968; Evans,, 1985) thus provides an explanation of the human capacity to understand novel sentences—we derive the meaning of novel constructions by recombining known components according to known syntactic rules.
@@
In fact, it can seem that tacit knowledge of a compositional grammar is the only adequate explanation of this capability. Chomsky argued that humans’ creative capacity is essentially unbounded (Chomsky,, 1957). To take a classic example, if we idealize away from constraints of memory and attention (as well as the finitude of a human lifetime!), a compositional subject could go on adding new conjuncts to the end of a well-understood sentence ad infinitum. Provided that they understand the meaning of the new conjunct (and conjunction itself), they will also understand these newly formed sentences. It is this open-endedness that makes an internalized compositional system seem like “the only game in town.” Alternative explanatory approaches of the time, like behaviorism, struggled to account for this form of unbounded extension. For Chomsky, Montague, Fodor, and others, the solution was clear: posit internal compositional mechanisms that support these behavioral capacities.
@@
This abductive inference to the best explanation forges a close link between creativity, productivity, and compositionality. Particular observations about productive or systematic behaviors in humans motivate the ascription of internal compositional mechanisms to ordinary speakers. Such arguments commit theorists to interpreting the compositionality principle as a psychological, rather than a metaphysical, thesis. [^6] To return to our analogy, if we aim to explain the behavior of someone solving a multiplication problem, what matters is how they determined the solution, in the subjective sense, rather than the metaphysical or objective grounds of this solution. This point deserves emphasis. To serve as a robust explanation of human language comprehension and observable linguistic behavior, these internal compositional mechanisms must be ‘psychologically real’ and actually employed as humans determine the meanings of novel expressions. [^7] This observation may seem obvious, but it is in fact easy to overlook. Even Davidson, one of the earliest and most influential proponents of compositional natural language semantics, fails to fully appreciate the point. For a detailed discussion of this tension in Davidson’s work, see Heck, (2013). For the proposed explanation to work, it is not enough that the language itself admits of a compositional analysis—after all, one and the same language may admit both compositional and non-compositional models.
@@
To see the point, consider a simplified ‘toy’ language (Evans,, 1985) with only 10 names (a, b, c…) and 10 predicates (F, G, H…), generating 100 possible sentences (Fa, Fb, Ga…). This language can be provided with a compositional model, along familiar lines. However, every admissible sentence in this language could also be generated by a very different, though ‘extensionally equivalent,’ grammar that has 100 different axioms specifying the meaning of each. How can we tell which of these models a speaker of this minimal language possesses? Many philosophers, from W. V. O. Quine, (1970) to Crispin Wright, (1981), were skeptical that such questions could be made empirically or behaviorally tractable. Evans, (1985), however, argues that there are clear empirical considerations that justify the ascription of the compositional system rather than the non-compositional one. In principle, we could scrutinize the speaker’s internal states and attempt to detect compositional structure. In practice, (and in keeping with the methodological assumptions of researchers in the Chomskyan tradition) we can instead rely on a wider class of behavior—compositional generalization behavior involving the introduction of novel names or predicate. If you introduce a new sentence, Fz, will the speaker be able to understand Gz and Hz? If so, this is evidence that they understand the language compositionally. Evans’ proposal about how to operationalize compositionality is very similar to those used in contemporary machine learning research (see section 3.1), and it demonstrates the central role that linguistic behavior plays in motivating cognitive explanation in terms of the compositionality principle.
@@
These theoretical developments close the gap between early philosophical theorizing about compositionality and the ongoing debates in contemporary cognitive science. Compositionality is understood as an explanatory psychological property, operationalized in terms of the specific kinds of productive and systematic generalization behaviors that humans exhibit (Evans,, 1985). In each case, the characteristic ‘behavioral signatures’ of compositionality are taken as evidence diagnostic of internalized compositional mechanisms, which serve to explain the behaviors in question. As Chomsky often emphasizes, this means that even the most abstract reaches of formal linguistics are properly understood as a branch of cognitive psychology (and ultimately biology, though developing an account of how symbolic systems are implemented in the brain is sometimes treated as a subsidiary enterprise; see Chomsky,, 1968).